OpenStudy (ttp):

A test consists of multiple choice questions worth 4 points and essay questions worth 8 points. The test has 22 questions that are worth a total of 100 points. Use a model to determine how many multiple choice questions are on the test. 14 16 19 22

OpenStudy (anonymous):

3

OpenStudy (anonymous):

8x +88-4x=100 4x+88=100 4x=22 x=3

OpenStudy (amistre64):

you can develop a system of 2 equations, with 2 unknowns

OpenStudy (ttp):

Isn't it 22

OpenStudy (amistre64):

m + e = 22 ; since there are 22 questions in all 4m+8e = 100 , 4 points for m and 8 ponts for e

OpenStudy (amistre64):

how do you solve systems of equations?

OpenStudy (ttp):

m+e = 22 4m + 8e = 100 use substitution e=22-m ->4m + 8(22-m) = 100 -4m = -76 m = 19 Am i right?

OpenStudy (amistre64):

let me chk, ill use elimination -8m -8e = -176 4m+8e = 100 ---------------- 4m = 76 i like it

OpenStudy (ttp):

Thank you! :)

OpenStudy (amistre64):

good luck ;)